OpenStudy (anonymous):

Help! Find the (perimeter) of a rectangular object which has a length of √128 feet and a width of √200ft. plz answer and give the final solution in simplified radical form.

OpenStudy (anonymous):

\[\sqrt{128} = \sqrt{64*2} = 8\sqrt{2}\] \[\sqrt{200} = \sqrt{100*2} = 10 \sqrt{2}\] perimeter = (length + width) * 2

OpenStudy (anonymous):